how do you solve this problem?
__8_ = 2
2x-4 = x

Found 3 solutions by richwmiller, stanbon, solver91311:
Answer by richwmiller(17219)

8/(2x-4)=2/x
8x=2*(2x-4)
8x=4x-8
4x=-8
x=-2

Answer by stanbon(75887)

------------------------
8/(2x-4) = 2/x
-------
8x = 4x-8
4x = -8
x = -2
============
Cheers,
Stan H.

Answer by solver91311(24713)



Cross-multiply. The product of the numerator on the left and the denominator on the right is equal to the product of the denominator on the left and the numerator on the right, that is:



Then collect terms and solve like any other single variable equation.

John